### v2

# PyTorchGeoNodes: Enabling Differentiable Shape Programs for 3D Shape Reconstruction

Abstract
We propose PyTorchGeoNodes, a differentiable module for reconstructing 3D objects and their parameters from images using interpretable shape programs. Unlike traditional CAD model retrieval, shape programs allow reasoning about semantic parameters, editing, and a low memory footprint. Despite their potential, shape programs for 3D scene understanding have been largely overlooked. Our key contribution is enabling gradient-based optimization by parsing shape programs, or more precisely procedural models designed in Blender, into efficient PyTorch code. While there are many possible applications of our PyTochGeoNodes, we show that a combination of PyTorchGeoNodes with genetic algorithm is a method of choice to optimize both discrete and continuous shape program parameters for 3D reconstruction and understanding of 3D object parameters. Our modular framework can be further integrated with other reconstruction algorithms, and we demonstrate one such integration to enable procedural Gaussian splatting. Our experiments on the ScanNet dataset show that our method achieves accurate reconstructions while enabling, until now, unseen level of 3D scene understanding.
Jointly recovered geometry and parameters of target objects from RGB-D scans
Refer to caption Refer to caption Refer to caption
Views of Target object Recovered Gaussians Procedural Gaussians editing
Figure 1: Our PyTorchGeoNodes is a framework that enables differentiable shape programs for 3D reconstruction. Our semantically-rich reconstructions of 3D objects from RGB-D scans enable understanding of important metric parameters of individual parts for different objects. These parameters are physically bound to the 3D reconstruction as PyTorchGeoNodes enables flow of gradients from the generated shape to the parameters of the object. Our PyTorchGeoNodes is modular, e.g. we can use it to enable procedural modeling for Gaussian splatting. In the example, our method recovers Gaussians for the target sofa and ’clones’ all details from the visible armrest to the occluded armrest in the scene. Our integration enables procedural editing of Gaussians and we can alter shape parameters to edit the reconstruction: e.g. by extending or shrinking width of the seat, removing backrest, left armrest, changing height of the backrest and depth of the seat etc.
1 Introduction
Various 3D object representations have already been considered for 3D scene understanding: voxels [31, 21], meshes [11, 38], SDFs [25], NeRFs [5], Gaussian Splatting [34], exemplars from a shape database [3]. SDFs, NeRFs, or Gaussian Splatting generalize to shapes unseen during training but produce artifacts for example when parts of the object are occluded in scenes captured “in the wild”. Their memory footprint can also become large. Using exemplars guarantees that the retrieved shapes look good as they are created by human designers, however the database may not contain shapes similar to the target object.
In this paper, we investigate the use of interpretable shape programs to represent objects. Shape programs are a form of procedural modeling [10] introduced in [35] for representing 3D objects. While shape programs have already been considered for 3D scene understanding [32, 27], their use in this context remains largely unexplored. As shown in Figure 1, our method can be used to accurately retrieve geometry and 3D parameters of target objects, together with its texture even in presence of partial occlusions.
Shape programs can be seen as computational graphs of parametrized geometric operations that generate various 3D shapes from an object category given input shape parameters, such as object dimensions as well as “semantic” properties like existence of armrest or backrest in case of chairs. Multiple similar parts, for example the legs of a chair or the shelves of a bookshelves—are obtained by first generating a single instance, duplicating it and applying geometric transformations such as rotations and translations as needed.
Thanks to this procedural modeling approach, shape programs generate shapes that remain appealing and that can fit many different instances. Alongside being interpretable, the shapes they generate are editable in an intuitive manner and have a very small memory footprint. In fact, they are already a popular tool for building reconstruction [39, 24, 19, 37] and the interest around shape programs for objects is growing [35, 15, 16, 27]. Blender [8] already provides support for them. A research challenge targeting reconstruction of walls, doors, and windows using such structured languages was organized by Meta in 2024 111[https://eval.ai/web/challenges/challenge-page/2115/overview (Accessed: November 2024)].
Recovering parameters of a shape program from an image is however challenging. Typically, a shape generated by a shape program depends on a number of parameters. Some parameters such as the width and height of the objects take continuous values but other parameters take only discrete values, e.g. the number of shelves in a bookshelf.
Optimizing on these parameters turns out to be tricky: For example, changing the number of shelves from 5 to 6 changes significantly the shape and appearance of a bookshelf. There have been a few attempts in the computer vision community to use procedural shapes [32, 27]. In [27] the authors train a neural network to predict the parameters of a shape program. While this method shows impressive results using synthetic data, we show in Section 5.3 that on real data such an approach fails to generalize well, due to the lack of real world image data annotated with shape parameters. Using synthetically generated data to train such models also performs very badly: While shape programs capture very well the global structure of objects, they do not represent details. As a result, the model can hardly generalize from synthetic images to real ones. In comparison, our proposed method utilizes an objective function which is not based on learning and does not need training data, hence it is able to generalize well to real world scenarios.
We introduce three main contributions which enable the estimation of shape programs’ parameters from images:
  • Our first contribution is a ‘compiler‘ that generates PyTorch [26] code for a shape program, which can be defined in Blender for example. This code generates 3D shapes that are differentiable with respect to the continuous parameters and can therefore be used easily with 3D reconstruction algorithms to optimize the continuous parameters.
  • Second, we contribute a method to obtain efficiently and reliably all shape program parameters from RGB-D scans: While our compiler can be used to optimize over the continuous parameters, the challenge of estimating the discrete parameters remains. We propose to adapt methods based on genetic algorithms to our problem of jointly optimizing the continuous and the discrete parameters of a shape program.
  • Our third contribution is an extension of shape programs to capture fine details and appearance of objects. We do this by augmenting our programs with Gaussians. Once we retrieved the shape parameters, we initialize our procedural Gaussians over the primitives and optimize them as in Gaussian splatting [17]. The key difference with regular Gaussian Splatting is that our Gaussians are constrained by the procedural model: The Gaussians sampled on ’cloned’ base primitives are tied together. e.g. as the legs of a chair are ’clones’ of the same base primitives, the occluded part of a leg can still be recovered if it is visible for another leg.
We evaluate our method on both synthetic and real scenes. In addition to the usual metrics, we also evaluate how well we reconstruct the shape program continuous and discrete parameters. Because existing datasets are not annotated with the value of these parameters, we manually annotated $176$ objects.
In summary, our main contribution is a general tool to create PyTorch code from Blender shape programs. In addition, we provide a method for estimating the parameters of shape programs from views together with the poses of the objects. Our PyTorchGeoNodes is a modular tool and it can be extended to novel use cases in 3D scene understanding. The source code is available to the general public.
2 Related Work
Since the field of retrieving 3D models for objects from images is very large, we focus here only on methods that can provide a light CAD model designed for specific categories: 3D model retrieval methods and Shape Programs.
2.1 3D Model Retrieval
3D model retrieval aims to retrieve a CAD model from a database of objects, such as the ShapeNet dataset [6], that maximizes some scoring function which is typically based on an input 3D representation or a single or multiple input images.
Some methods approach CAD model retrieval as a classification problem [3, 23]. For a given input a neural network can be trained to predict a probability distribution over all models in the database. However, these direct approaches do not scale well to large object databases such as the ShapeNet database. Another popular approach is computing a similarity metric between synthetic objects and real inputs. There are several possible ways of computing such metrics. It is possible to render an image from a synthetic object, and then for a synthetic-real image pair, we can compute and compare image descriptors [2, 13, 22] or encode images into embedding vector representations and calculate the similarity in the embedding space [20].
While such approaches have displayed promising results, one clear limitation is expressiveness of the database of objects. In fact, even with large databases of objects, in practice we often encounter objects in real scenes that are not represented in the database, or we might encounter modified or articulated objects. For example, we can modify seat height, arm rest and back support position of a work chair but the object in the database appears in the default setting which will introduce errors during retrieval. While it is possible to perform mesh refinement and deformation [4, 33, 14], these approaches typically introduce artifacts and deformations in the reconstruction and the final object shape might not be representative of the actual object category after refinement.
2.2 Shape Programs
Procedural modelling is an attractive way of generating 3D shapes in computer graphics, and as such it is an interesting feature of different 3D modelling software. In general, procedural modelling is a family of techniques that generate 3D content by abiding to a set of rules. A variety of methods were proposed that combine procedural modeling with machine learning approaches, summarized in detail in [30]. For example, [35] introduced a procedural language that enables the generation of a variety of shapes in form of so called shape programs. In [35], shape programs are written in Domain Specific Language (DSL) for shapes with two types of statements. “Draw” statements describes shape primitive, such as cuboid or a cylinder, together with its geometric and semantic properties. “For” statements represent loops of sub-programs including parameters that indicate how the sub-program should be repeatedly executed.
[35] also showed that it is possible to learn to infer shape programs that describe input 3D shapes. However, as it is based on basic geometric primitives, it becomes difficult to model more complex geometries, and resulting shapes can be non-representative of the target category. Other methods suffer from the same limitations [15, 16].
GeoCode [27] recently introduced hand-designed shape programs as a mapping between 3D shapes of a specific object category and human-interpretable parameter space. The language used for developing shape programs is based on the geometry node feature of Blender, an open-source modeling software, which, in contrast to [35], allows designing of more expressive shape programs that can represent and interpolate between a variety of shapes by simply modifying human-interpretable program’s parameters.
Figure 2: A procedural model in Blender generating dividing boards of a cabinet. The computational graph is designed and visualized in Blender using Geometry Nodes feature. Underneath, we show how we abstract the different nodes using PyTorch tensors and PyTorch3D meshes. The input node takes input parameters, here {Width: 0.5, Dividing Board Thickness: 0.04, Height: 0.6, Number of Dividing Boards: 5, Board Thickness: 0.04} and feeds them to a series of operations. The blue nodes are arithmetic and concatenation nodes, which transform input parameters and feed the results to geometry nodes, in green. In this example, we generate a cuboid mesh and instantiate a line of points which generates the final geometry for dividing boards. In practice, this shape program is part of a larger shape program for modeling cabinets.
Some nodes define a shape primitive or pre-designed more complex shapes, e.g., chair seat or chair leg. Other nodes represent geometric transformations on individual shapes, or how two shapes should be connected, e.g., legs should always be connected to the bottom of the seat. There are nodes representing mathematical operations such as addition and multiplication, and there are nodes that are used to replicate specific shapes, e.g., in case of chair legs we can add a replicator node to create four instances of a chair leg of the same shape. There are also nodes that serve as control statements, e.g., it is possible to switch between a four-legged chair and a one-legged chair. A pre-defined set of input program’s parameters allow non-expert user to easily control shape-generation process. [27] also showed that it is possible to learn the mapping from input point clouds of 3D shapes, or a sketch, to program’s parameters.
3 PyTorchGeoNodes
In this section, we discuss shape programs and how they can be beneficial for tasks in 3D scene understanding. We then describe our framework for enabling efficient gradient-based optimization for shape programs and, in addition, allowing to translate shape programs designed in Blender to PyTorch code to enable differentiable optimization. In Section 4, we explain how our framework can be applied for reconstructing 3D objects from images.
We consider shape programs as defined in Blender, i.e., as a computational graph that computes 3D shapes given a number of input parameters. Such graphs contain nodes of different types; in our framework, we currently consider the following node types:
  • An ’Input’ node contains input parameters of the shape programs and their value ranges. We consider floating, integer, and boolean values;
  • A ’Math’ node performs mathematical operations on its input (addition, subtraction, multiplication, division, and comparison);
  • A ’Switch’ node switches between two inputs conditioned on a boolean input;
  • A ’Combine’ node concatenates three input values into a 3D vector;
  • A ’Primitive’ node creates a geometric primitive, e.g., a cuboid or a cylinder;
  • A ’Transform’ node applies scaling, rotation, and translation to the input geometry;
  • ’Mesh Line’ node creates a tensor of 3D points on a line from input vectors that define the start location and the offset of the line, and an integer that defines the number of points to be generated;
  • A ’Points on instances’ node clones input geometry at given input 3D points;
  • A ’Join geometry’ node merges geometries;
  • An ’Output geometry node’ generates the final mesh for the shape.
Note that for example, the Transform node depends on continuous, the Points on instances node on integer, and the Switch node on boolean.
Our framework provides differentiable computational nodes that reimplement the functionalities of the individual geometry nodes in Blender as we show in Figure 2. More exactly, for every node type in Blender, we implement a corresponding node type with the same functionalities using PyTorch, or PyTorch3D [29] in case of geometric operations. In practice, we designed shape programs directly using the Blender interface. Given a shape program created in Blender, our parser generates a differentiable PyTorchGeoNodes code. We can then use this code in PyTorch in an optimization procedure where we fit shape parameters to generate a shape consistent with the input scene.
In the supplementary material, we detail our design choices and technical details which make our implementation computationally efficient.
Figure 3: Fitting shape parameters with PyTorchGeoNodes. At every iteration, our genetic algorithm generates indiviudals of shape parameters and object poses. We parse 3D shapes from individual parameters using PyTorchGeoNodes and perform shape parameter optimization based on loss that evaluates how well the shape fits the input scene. In the next iteration, the optimized set of individuals is used to update the pool of individuals based on selection of the fittest principle.
4 Search Algorithm
We leverage our PyTorchGeoNodes framework to retrieve shape parameters and geometry of a target object in an RGB-D scan. We look for shape parameters ${\cal P}^{*}$ by minimizing an objective function that encourages the shape to fit the point cloud reconstructed from the images while lying on the floor:
${\cal L}({\cal P})=$ $\lambda_{\text{CD}}{\cal L}_{\text{CD}}({\text{PC}},{\text{Sh}}({\cal P}))+\lambda_{\text{G}}{\cal L}_{\text{G}}({\text{G}},{\text{Sh}}({\cal P}))+$
$\lambda_{\text{Fl}}{\cal L}_{\text{Fl}}({\text{PC}},{\text{Sh}}({\cal P})))\>.$ (1)
${\cal L}_{\text{CD}}({\text{PC}},{\text{Sh}}({\cal P}))$ measures the single-direction chamfer distance between PC, the point cloud reconstructed from the depth maps and object masks, and ${\text{Sh}}({\cal P})$, the shape generated by the shape program. We follow pre-processing from [1] which relies on sparse point cloud instance segmentation of the object and use a segmentation model [18] on images to obtain the object masks. We then back-project these masks to obtain a dense target point cloud. ${\cal L}_{\text{G}}({\text{G}},{\text{Sh}}({\cal P}))$ penalizes occlusions in the scene caused by ${\text{Sh}}({\cal P})$. G is an occlusion grid obtained by densely sampling 3D points along rays between the camera and surfaces for different views:
${\cal L}_{\text{G}}({\text{G}},{\text{Sh}}({\cal P}))=\sum_{p\in{\text{Sh}}({\cal P})}\begin{cases}\stackrel{{\scriptstyle\rightarrow}}{{{\text{CD}}}}(p,{\text{PC}})&\text{if }{\text{G}}(p)\\ 0&\text{otherwise},\end{cases}$ (2)
where $\stackrel{{\scriptstyle\rightarrow}}{{{\text{CD}}}}(p,{\text{PC}})$ is the distance of point $p$ to PC . ${\text{G}}(p)=\stackrel{{\scriptstyle\rightarrow}}{{{\text{CD}}}}(p,{\text{PC}})<\omega$ determines whether $p$ is on grid G. We set threshold $\omega=5cm$ in our experiments. Since G is generated using input views for the scene, this loss does not penalize points that are occluded in the scene, similar to render-and-compare objectives. The term ${\cal L}_{\text{Fl}}$ simply enforces the objects to stay on the floor plane:
${\cal L}_{\text{Fl}}({\text{PC}},{\text{Sh}}({\cal P})))=|{\text{Fl}}({\text{PC}})-{\text{Fl}}({\text{Sh}}({\cal P}))|_{2}\>,$ (3)
[…]
We use our PyTorchGeoNodes framework to compute ${\text{Sh}}({\cal P})$. Thanks to this framework, PyTorch can compute the derivatives of ${\text{Sh}}({\cal P})$ and thus the derivatives of ${\cal L}$ with respect to the continuous parameters in ${\cal P}$. However, some parameters in ${\cal P}$ are discrete, i.e., integers or booleans. Therefore, we cannot optimize ${\cal L}$ using a simple gradient descent. Also, ${\cal L}$ has no special form that we can exploit to optimize it.
To solve this problem, we propose to resort to genetic algorithms to optimize the parameters and combine it with gradient descent to further refine the continuous parameters.
4.1 Shape Parameter Search
A genetic algorithm is a stochastic optimization algorithm that iteratively applies random manipulations of high-dimensional search parameters to identify the most promising complex search subspaces, subject to an objective function. In contrast to gradient-based techniques, it can optimize discrete and continuous values and the objective function can be very general and non-differentiable.
The algorithm starts by randomly generating instances of parameter values ${\cal P}$, individuals, that form an initial set called population of search parameters. Every iteration has three phases: crossover and mutation for exploring search subspaces, and selection for exploiting the most promising subspaces. Crossover randomly selects pairs of individuals, parents, and mixes them by combining the individual parameters to create a new offspring. Therefore, an offspring lies in the search subspace at intersection of parents’ search spaces. Mutation then randomly alters offsprings’ parameters to further explore this subspace. Selection evaluates the quality of the population, and based on survival of the fittest strategy, it takes the best performing individuals to form the population for the next iteration. Further details and hyper-parameters are given in the supplementary material.
Our adaptation of genetic algorithms to our problem is described in details in the supplementary material. The main points are given below. We generate an initial population by randomly generating instances of all program’s parameters ${\cal P}$ representing individuals in the population. We draw random values from the empirical distributions of the parameters. The crossover operation generates offspring $P_{1,2}$ from a pair of parents ${\cal P}_{1}$, ${\cal P}_{2}$. Mutation varies depending on the parameter type. For discrete parameter $d\in{\cal P}_{D}$, we adapt its value $v_{d}$:
$v_{d}\leftarrow\begin{cases}v_{s}\in_{R}{\cal V}_{d}&\quad\text{with probability }P_{m}\\ v_{d}&\quad\text{otherwise},\end{cases}$ (4)
where operator $\in_{R}$ randomly selects value $v$ from the set of valid discrete values ${\cal V}_{p}$. $P_{m}$ is a hyper-parameter that defines the probability of mutating parameter values. In case of a continuous parameter $c\in{\cal P}_{C}$, we adapt its value $v_{c}$ as:
$v_{c}\leftarrow\begin{cases}v_{s}\in_{R}{\cal V}_{c}+{\cal N}(0,\sigma)&\text{with prob. }P_{m}\cdot P_{m}\\ v_{c}+{\cal N}(0,\sigma)&\text{with prob. }P_{m}\\ v_{c}&\text{otherwise},\end{cases}$ (5)
where ${\cal V}_{c}$ is a discretisation of values for parameter $c$ that we further discuss in the supplementary. ${\cal N}(0,\sigma)$ is normal distribution with mean value set to 00 and standard deviation $\sigma$. Finally, we evaluate offsprings, append them to the current population and we select the best individuals based on a score function, in our case negative of objective in Equation 1, to form the population for the next iteration.
Optimization step. Using PyTorchGeoNodes, during the selection phase of every iteration, we perform a gradient-based optimization to further refine values of continuous shape parameters of individuals before selecting the population for next iteration. After a specified number of iterations the search concludes and we return shape parameters of the best performing individual.
4.2 Integrating Gaussians Splats
Shape programs explicitly capture symmetries and similarities between repeated structures, while this aspect is lacking in many representations for 3D reconstruction. On the other hand, earlier shape programs do not capture well details and appearance.
In this section, we show we enable procedural Gaussian splats by integrating them within shape programs and our PyTorchGeoNodes framework. This allows us to capture fine details of objects while keeping the advantages of shape programs. First, we explain how we augment the primitives with ’Gaussian primitives’ while keeping the graph of PyTorchGeoNodes differentiable, and then, we introduce our render-and-compare objective function.
Refer to caption Refer to caption
Original Mesh Augmented with Gaussians
Figure 4: We integrate Gaussians directly into shape programs and PyTorchGeoNodes, to have the advantages of both procedural models and Gaussian splatting.
Extending PyTorchGeoNodes with Gaussian primitives. To enable integration of our procedural graph to support Gaussian splats, we extend the nodes creating mesh primitives to also create trainable Gaussian parameters. We subsample the primitives meshes into many triangles and create Gaussians from these triangles. We initialize the Gaussians means, scales, and quaternions to fit the mesh triangles as we show in Figure 4. Their colors are initialized based on the closest 3D point in the scan. Their opacities are fixed to a high value. As with other primitives, the graph transforms the Gaussians to form a complete set of Gaussians, some of them sharing parameters due to the cloning in the graph.
Optimizing the Gaussians. We noticed that it is more robust to, rather than jointly, first estimate the shape parameters and then optimize the Gaussians. Given a set of optimal shape parameters ${\cal P}$, we minimize the render-and-compare loss as in other Gaussian splatting works, by rendering the Gaussians ${\cal G}({\cal P})$ into scene views [40, 17, 12]:
${\cal L}_{\text{RGB}}+\lambda_{1}{\cal L}_{\text{D}}+\lambda_{2}{\cal L}_{\cal P}+{\cal L}_{\text{reg}}\>,$ (6)
where ${\cal L}_{\text{RGB}}$ and ${\cal L}_{\text{D}}$ are standard losses on color and depth consistency used in Gaussian splatting literature [40, 7, 36].
${\cal L}_{{\cal P}}={\text{CD}}({\text{Sh}}({\cal P}),{\cal G}({\cal P}))$ measures the consistency between the shape parameters and the Gaussians by relying on the chamfer distance. ${\cal L}_{\text{reg}}$ is a regularizer term:
${\cal L}_{\text{reg}}=$ $\lambda_{3}|M({\cal G}({\cal P}))|_{2}+\lambda_{4}|S({\cal G}({\cal P}))|_{2}+$ (7)
$\lambda_{5}\sum_{G\in{\cal G}({\cal P})}\text{Var}(C(G)|,$ (8)
to enforce mean offsets $M({\cal G}({\cal P}))$ and scales offsets $S({\cal G}({\cal P}))$ of Gaussians to be small and the colors of the Gaussians on the same primitive $C(G)$ to remain close to each other. We set $\lambda_{1}=0.1,\lambda_{2}=0.1,\lambda_{3}=100,\lambda_{4}=10.00,\lambda_{5}=1$ in our experiments. In addition, we observed that optimizing object Gaussians and background Gaussians jointly at the same time, with loss for background locations set to $\lambda_{bgd}({\cal L}_{\text{RGB}}+\lambda_{1}{\cal L}_{\text{D}})$ and $\lambda_{bgd}=0.5$ ensures that object Gaussians remain consistent around edges.
5 Validation
We validate our PyTorchGeoNodes framework and our optimization method on three common furniture categories: Sofa, with $18$ parameters, Chair with $20$, and Table, with $10$, for which we created shape programs in Blender. We designed our shape programs to capture interpretable object structures that are relevant for objects in ScanNet, allowing unseen-before reasoning about high- level object parameters. More details are provided in the supplementary material.
Our method jointly estimates the pose and shape parameters of the object. The orientation is defined as the rotation around the up-axis as done in to [1], which is sufficient for indoor scenes.
We validate our method on object shape reconstruction for real world scenes from the ScanNet dataset [9]. ScanNet is a challenging dataset of indoor environment RGB-D scans: Furniture is only partially visible and highly occluded, and depth maps are incomplete due to missing sensor values—this happens for dark and reflective materials, for example. We manually annotated $176$ objects for Sofa, Chair, and Table. We also provide additional results on automatically generated synthetic scenes in the supplementary material including an additional object category.
5.1 Evaluation Criteria
Parameter Model from [27] Model from [27] CD CD Genetic Genetic
[…]
Continuous Parameters Mean Absolute Difference to Ground Truth for Table (↓)
Width [cm] 0.74 0.22 0.21 0.21 0.15 0.08
Height [cm] 0.34 0.14 0.15 0.15 0.12 0.06
Depth [cm] 0.4 0.2 0.22 0.22 0.13 0.08
Mid Leg X Scale 0.19 0.13 0.16 0.14 0.1 0.07
Mid Leg Y Scale 0.16 0.14 0.11 0.11 0.08 0.08
Legs Offset Y 0.35 0.34 0.37 0.37 0.31 0.27
Legs Offset X 0.32 0.32 0.36 0.36 0.15 0.13
Legs Scale Y 0.17 0.15 0.18 0.19 0.13 0.1
Mid Board Z Scale 0.21 0.18 0.18 0.18 0.17 0.14
Cabinet Width Scale 0.15 0.1 0.15 0.11 0.1 0.05
Top Thickness [cm] 0.03 0.03 0.02 0.02 0.02 0.02
Legs Size [cm] 0.03 0.03 0.03 0.03 0.01 0.01
Discrete P. Classification Accuracy for Table (↑)
Top Shape 0.78 0.78 0.72 0.72 0.94 0.94
Legs Type 0.39 0.39 0.6 0.58 0.78 0.78
Has Cabinet Leg 0.71 0.71 0.57 0.55 0.78 0.86
Has Mid Board 0.71 0.71 0.55 0.49 0.76 0.76
Reconstruction quality across all categories (↓)
Chamfer distance [m] 0.52 0.21 0.16 0.16 0.14 0.11
Abs. rotation error [rad] 0.45 0.12 0.13 0.11 0.08 0.06
Table 1: Quantitative results of shape program parameters between different baselines for category Table, and general reconstruction metrics across all categories. We observe that GeoCode [27] performs worst due to the domain gap, but the optimization helps improve the performance. Coordinate Descent (CD) does better but is prone to local minima. Genetic algorithm outperforms other baselines by a significant margin.
Evaluation of shape parameters. Each object category has properties which are an important aspect of the reconstruction. This is neglected by current state-of-the-art methods. Our approach integrates these properties as shape parameters into the reconstruction and we evaluate them by comparing the reconstructed and ground truth parameters of the target objects. For the experiments with ScanNet, we manually annotated all relevant objects with shape parameters.
Data pre-processing. We assume that for the target object in the RGB-D scan, we have an initial estimate for the 3D center of the target object, here we follow preprocessing from [1]. Then, we estimate 2D segmentation masks of the object used to generate dense point clouds in the objective function in Equation (1).
5.2 Search Baseline Details
In our evaluations of our search algorithm, we compare results of our genetic algorithm to two baselines:
  • GeoCode [27] introduces a supervised learning method based on an encoder-decoder architecture that takes a target point cloud as input and outputs a value for target program parameters. We use DGCNN [28] as encoder, and the resulting embedding of size $128$ is shared between separate decoders that predict values for program parameters. More precisely, we use three feed-forward layers per parameter, each of size $256$. For discrete parameters, we add a classification head as the output layer with number of outputs matching the number of different discrete values. For continuous parameters, we simply regress the output value. Finally, we perform gradient descent that further refines predictions by passing predicted parameter values to PyTorchGeoNodes and optimizing the objective function. We train it by randomly generating $10000$ instances of ground truth shape program parameters and rotations and then generating synthetic point clouds from these parameters using PyTorchGeoNodes.
  • Our implementation of coordinate descent iteratively optimizes individual parameters of the shape program. First, we initialize the estimate of parameters randomly. Then, at every iteration, we take the current estimate and replace value of the selected parameter by all valid discrete values and calculate the objective term. In case of continuous values we first discretize them as we explain in the supplementary. In case this improves the objective, we update the current estimate. Finally, we perform gradient descent that further refines the current estimate by passing it to PyTorchGeoNodes and optimizing the objective function. We repeat the process for all different parameters and loop 3 times which approximately matches the execution time of our genetic algorithm.
5.3 Validation on the ScanNet Dataset
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Shape from recovered object parameters aligns very well with the target object also when projected into sampled images of the scene. In the first row, we reconstruct a variety of chairs along with many detailed parameters including seat thickness, thickness and offsets of side and front/back support legs, existence of different semantic elements like ’star-shaped’ legs and its exact rotation. In the second row, we correctly estimate shape of table top and its measurements, including width, depth and thickness, parameters of table legs, and existence of middle shelf including its offset from the bottom. In the third row we reconstruct parameters of ’L-element’ of sofa including its depth and width, existence and size of legs, and existence of individual armrests, backrest, and the corresponding measurements. We show more results in the supplementary material.
We evaluate our approach on challenging real-world scenarios on validation scenes from ScanNet [9], which provides RGB-D scans of indoor scenes. We selected a subset of object instances from the Scannotate dataset [1] for which we manually annotated 57 sofas, 52 chairs, and 67 tables.
Table 1 evaluates the quality for object parameters reconstruction for the ’table’ category and general reconstruction quality across different categories. We report the numbers for ’chair’ and ’sofa’ in the supplementary material. We measure the classification accuracy for discrete parameteres and absolute difference for continuous parameters.
We observe that the supervised learning method from [27], trained on synthetic data, generalises poorly to the scene from ScanNet, while coordinate descent baseline is prone to local minima. Our genetic algorithm performs very well across all metrics except for parameters defining very small details. Visually, the models recovered by our method are comparable in quality to the ground truth and we confirm the benefits of our refinement procedure. Figure 5 and Figure 6 show some qualitative results. In the supplementary we show additional results, and include a downstream application for point cloud segmentation based on our reconstructions.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
       (a) (b) (c)
Figure 6: (a) We overlay shape reconstructed using our retrieved shape parameters with the image and (b) compare it to the ground truth. (c) Our integration of Gaussians into PyTorchGeoNodes enables appealing 3D reconstructions.
6 Conclusion
Our code is publicly available and we hope this will encourage other researchers to develop the field and further extend possibilities with PyTorchGeoNodes. While procedural models / shape programs are an attractive representation of objects, there are indeed still many things to develop: e.g. extensions to consider deformable and articulated objects and to integrate textures, or extensions for scene-level procedural models / programs (learned or designed in Blender for example) that would enable reasoning about object relationships and scene physics. Modularity of PyTorchGeoNodes enables integrations with other types of 3D representation, as we demonstrate with our integration of Gaussian Splatting. Moreover, differentiable rendering of PyTorchGeoNodes enables development of deep learning methods for creating shape programs from images in a self-supervised way. Finally, a major limitation is that our shape programs need to be designed by humans for different categories which can lead to limited expressiveness when applying our approach in the wild. Learning shape program designs from database of shapes would make our method easily applicable to novel categories.
Acknowledgement
This project was funded by the European Union (ERC Advanced Grant explorer Funding ID #101097259). This work was performed using HPC resources from GENCI-IDRIS (Grant 2024-AD010615585).